What is prime factorization? How do I find the prime factorization of 90?
Answer by xcentaur(357)

Prime numbers are numbers which cannot be divided by another other number completely.example:2,3,5,7,11,etc


Factors are numbers which divide other numbers without leaving a remainder
eg: 3 is a factor of 9,but 2 is not a factor of 9 since it leaves a remainder of 1.


Now prime factors means all the prime numbers that are factors of the given number


we have been given 90
90 can be divided by short division as,

2|90
3|45
3|15
5| 5
 | 1
----

So the prime factors of 90 are:
90=2*3*3*5


this means suppose you have a number,say 28,28 can be divided by 14.but 14 is not a prime number,14=7*2.so we can write 28=2*7*2.those are its prime factors


Hope this helps,
Prabhat